OpenStudy (anonymous):

a rectangular plot of ground is 30 meters longer than it is wide the area is 15000 which eqaution will help you find the dimensions

OpenStudy (campbell_st):

let the width be w then the length is w + 30 so the area is \[1500 = w \times (w + 30)\] which means you have a quadratic equation \[w^2 + 30w -1500 = 0\] solve the equation for w, and only use the positive solution. then you can find the length and width. hope it helps

OpenStudy (anonymous):

yes i got it right thank you